# Energy-Efficient Channel Decoding for Wireless Federated Learning: Convergence Analysis and Adaptive Design

### v2

@@ III-C Decoding Iteration Control
$Q_{r}=\frac{1}{\chi_{0}}\mathcal{O}\left(2\log\ r\right)+\frac{1}{\chi_{0}}\mathcal{O}\left(\frac{1}{\epsilon}\log\frac{1}{\epsilon}\right),$ (17)
which provides a guideline to select the maximum number of LDPC decoding iterations in the $r$-th communication round. Specifically, $Q_{r}$ can be set as $\alpha\log r+\beta$, where $\alpha$ and $\beta$ are constants. However, (17) cannot be directly utilized because $\alpha$ and $\beta$ are code-specific and depend on the receive SNR as well as channel decoding algorithms. Thus, we propose a practical approach by the mapping relationship which captures the relationship between BER and the maximum number of decoding iterations at different levels of receive SNR. Entries in the mapping relationship can be obtained through empirical measurements specific to the LDPC code used in the wireless FL system. Fig. 5 shows the exampled simulation results for a rate 1/2 LDPC code with a block length of 1008 bits and binary phase shift keying (BPSK) modulation over an additional white Gaussian noise (AWGN) channel with $\text{SNR}=1.5$ and 2.5 dB. Data in this figure can be stored in memory [43]. From the mapping relationship, we can get the corresponding maximum number of decoding iteration for a target BER for clients with different receive SNR.
Figure 5: The relation between BER and LDPC maximum decoding iterations under different SNRs.
The wireless FL procedures with adaptive LDPC decoding are summarized in Algorithm 1. The algorithm starts with an initial target BER, denoted as $b_{0}$, which is set to a relatively high value, i.e., $10^{-1}$ in our experiments. The target BER at the last communication round, denoted as $b_{R-1}$, is set as a small value, i.e., $10^{-4}$ in our experiments. The concrete expression of $b_{r}$ in our experiment is $b_{r}=\frac{0.0999R^{2}}{(R^{2}-1)(r+1)^{2}}+\frac{10^{-4}R^{2}-0.1}{R^{2}-1}$. For each communication round, the BER is scaled according to (11), and the corresponding maximum number of LDPC decoding iterations in that round is obtained from the mapping relationship.
@@
Note that the number of entries in the mapping relationship depends on the possible discrete BER and SNR values. Since this number is not large, the memory overhead for storing the mapping relationship is expected to be small. Additionally, reading data from the memory takes constant time and does not introduce additional LDPC decoding latency.
@@ IV Experimental Results
In this section, we validate the effectiveness of the proposed adaptive LDPC decoding algorithm via numerical experiments. First, we introduce the experimental settings. Then, the performance of the proposed algorithm, including test accuracy and energy consumption, is compared with a baseline with a fixed number of LDPC decoding iterations.
Figure 6: Fixed decoding schemes: On the client side, setting the LDPC maximum decoding iterations to 6, 12, 18, 24, and 52 can achieve a BER of around $10^{-1}$, $10^{-2}$, $10^{-3}$, $10^{-4}$, and $10^{-5}$, respectively.
@@ IV-A Experimental Settings
(a) Experiment of LeNet-300-100 on MNIST.
(b) Experiment of vanilla-CNN on Fashion-MNIST.
(c) Experiment of CNN on CIFAR-10.
(d) Experiment of vanilla-CNN on non-IID Fashion-MNIST.
Figure 7: Experiments on different datasets showing how fixed BER affects FL accuracy.
(a) Consumed decoding iteration for each frame.
(b) The distributed for different consumed iterations.
Figure 8: Example of vanilla-CNN on Fashion-MNIST. The 8-bit quantized vanilla CNN model goes through AWGN channel with $E_{b}/N_{0}$ of 2.5dB, and is decoded by LDPC of block-length of 1008 bit.
@@
System Setups: We consider an FL system with 10 clients. All clients are assumed to be selected in every communication round and they communicate with the server using rate 1/2 min-sum LDPC code with a code length of 1008 bits[44]. Besides, binary phase shift keying (BPSK) is adopted as the modulation scheme. In each communication round, the number of local SGD steps is set as 5 with the learning rate and batch size as 0.01 and 64, respectively. The receive SNR of all the wireless channels is assumed to be 2.5 dB.
@@
Datasets: We conduct experiments on three dataset, including MNIST[45], Fashion-MNIST[46], and CIFAR-10[47]. The training datasets are distributed among clients in an IID manner. We further consider one experiment where the Fashion-MNIST data are non-IID among clients. In particular, each client was assigned with two categories of training data, similar to the setup in [24]. The test datasets are used for validation at the server.
@@
DNN Models: For the MNIST dataset, the LeNet-300-100 model[48] is employed, which is a fully connected neural network consisting of two hidden layers and one softmax layer, with approximately 0.27 million parameters. For the Fashion-MNIST dataset, a vanilla CNN [3] with two 5×5 convolution layers, one fully connected layer, and one softmax layer is employed, comprising around 0.58 million parameters. For the CIFAR-10 dataset, a 7-layer CNN model with four 3×3 convolution layers, two fully connected layers, and one softmax layer is employed, which contains approximately 9.07 million parameters.
IV-B Learning Performance of Conventional Schemes
(a) Max.iter.
(b) Learning curves.
(c) Average iterations.
(d) BER.
(e) Final learning accuracy.
(f) Total decoding energy.
@@ IV-B Learning Performance of Conventional Schemes
Figure 9: Experiment for LeNet-300-100 on MNIST dataset. Compared with the conventional fixed schemes, our method that adaptively increases the maximum decoding iteration to 24 can achieve the same final accuracy of 98%, and reduce the decoding energy by 6.7mJ.
(a) Max.iter.
(b) Learning curves.
(c) Average iterations.
(d) BER.
(e) Final learning accuracy.
(f) Total decoding energy.
@@
Figure 10: Experiment for vanilla CNN on IID Fashion-MNIST dataset. Compared with the conventional fixed schemes, our method that adaptively increases the maximum decoding iteration to 24 can achieve a similar final accuracy of 91%, and reduce the decoding energy by 14.1mJ.
@@
In this subsection, experimental results for the baseline scheme are presented. As shown in Fig. 6, it leads to approximately constant BERs over the training process.
Next, we investigate the impact of a constant BER on the performance of wireless FL. Fig. 7 summarizes the test accuracy with different maximum numbers of decoding iterations. It can be observed that for all the four FL tasks, when the BER is higher than $10^{-3}$, the learning performance was significantly impaired. Lowering the BER to $10^{-4}$ leads to improved accuracy and faster convergence towards wireless FL with error-free communication. However, further decreasing the BER to $10^{-5}$ does not yield any significant improvement in test accuracy.
@@ IV-C Learning Performance of the Proposed Scheme
(a) Max.iter.
(b) Learning curves.
(c) Average iterations.
(d) BER.
(e) Final learning accuracy.
(f) Total decoding energy.
@@
Figure 11: Experiment for 7-layer CNN on CIFAR-10 dataset. Compared with the conventional fixed schemes, our method that adaptively increases the maximum decoding iteration to 24 can achieve a similar final accuracy of 77.6%, and reduce the decoding energy by 190mJ.
(a) Max.iter.
(b) Learning curves.
(c) Average iterations.
(d) BER.
(e) Final learning accuracy.
(f) Total decoding energy.
@@
Figure 12: Experiment for vanilla CNN on non-IID Fashion-MNIST dataset. Compared with the conventional fixed schemes, our method that adaptively increases the maximum decoding iteration to 24 can achieve a similar final accuracy of 54.1%, and reduce the decoding energy by 13.9mJ.
@@
In this subsection, the proposed adaptive LDPC decoding scheme is compared to the baseline scheme. Specifically, the learning process of the proposed scheme is assigned with an initial BER of $10^{-1}$ and a final BER value of $10^{-4}$. The desired BER at every communication round is calculated via (11) and the maximum number of LDPC decoding iterations is determined according to the mapping with BER as stated in Section III. In the experiments, the initial and final maximum numbers of LDPC decoding iterations are thus set to 6 and 24, respectively. For the baseline scheme, a maximum of 24 LDPC decoding iterations are used in each communication round, which corresponds to a BER of $10^{-4}$.
@@
Fig. 8(a) shows an example on the Fashion-MNIST (IID) datasets to illustrate the actual number of decoding iterations for each frame of data bits. It can observed that only a few frames exhaust all the decoding iterations which is set to a large value of 52. Note that the decoding energy consumption is proportional to the actual number of decoding iterations used for the FL tasks. From the distribution on the actual number of decoding iterations shown in Fig. 8(b), it is observed that 18 decoding iterations are sufficient for most cases, implying the significant reduction of energy consumption.
Fig. 9 summarizes the experimental results for LeNet-300-100 on the MNIST (IID) dataset. Specifically, Fig. 9(a) shows how the maximum number of LDPC decoding iterations varies with communication rounds in the proposed scheme. While the baseline scheme with a fixed number of maximum decoding iterations results in a constant BER, the proposed scheme achieves the desired BER by adopting a linearly increasing maximum number of decoding iterations. Fig. 9(b) shows the test accuracy achieved by different schemes. It is observed that setting the maximum number of decoding iterations to 6 (which corresponds to a BER of $\sim 10^{-1}$) severely affects the learning accuracy. Similarly, when the maximum number of decoding iterations is set to 12 (which corresponds to a BER of $\sim 10^{-2}$), the training process still fails to converge. By setting the maximum number of decoding iterations as 18 (which corresponds to a BER of $\sim 10^{-3}$), the training process converges, but with a noticeable loss in test accuracy compared to the case when the maximum number of decoding iterations is 24 (which corresponds to a BER of $\sim 10^{-4}$). However, further increasing the maximum number of decoding iterations to 52 does not lead to much accuracy improvement.
@@
The proposed scheme, which linearly increases the maximum number of decoding iterations from 6 to 24, achieves the same test accuracy as those with fixed 24 or 52 decoding iterations. Fig. 9(c) illustrates how the average number of decoding iterations over all DNN parameters varies during the training process. The proposed scheme can reduce the average number of decoding iterations and thus the overall energy consumption. Fig. 9(d) depicts the BER of the proposed scheme over communication rounds, which scales as $\mathcal{O}(1/(r+1)^{2})$ from $10^{-1}$ to $10^{-4}$. Fig. 9(e) shows the final test accuracy of different schemes at the 50-th communication round. It can be observed that the proposed scheme achieves the same accuracy as the baseline schemes that always maintain a very low BER of $10^{-4}$ and $10^{-5}$ throughout the training process. Fig. 9(f) illustrates the decoding energy consumption on clients under different schemes. It is obtained by multiplying the total number of LDPC decoding iterations with the LDPC decoding energy efficiency reported in [29]. Compared to the baseline scheme with a maximum of 24 decoding iterations, the proposed scheme reduces the overall decoding energy consumption by 19.4%.
@@
TABLE III: Experimental Results of Different Schemes. Our scheme can achieve similar accuracy while saving around 20% of the decoding energy compared to the best-fixed scheme.
Final learning accuracy
@@
MNIST 96.74% 97.96% 98.06% 98.02%
Fashion-MNIST 86.18% 90.63% 91.24% 91.16%
CIFAR-10 76.72% 77.38% 77.57% 77.68%
non-IID Fashion 51.3% 54.1% 54.3% 54.2%
Total consumed decoding energy (mJ)
@@
MNIST 33.87 34.51 34.68 27.81
Fashion-MNIST 74.23 76.10 76.29 62.17
CIFAR-10 925.10 948.44 951.36 757.80
non-IID Fashion 74.22 76.11 76.27 62.38
Figs. 10, 11, and 12 present the experimental results for Fashion-MNIST (IID), CIFAR-10 (IID), and Fashion-MNIST (non-IID) datasets, respectively. The results are similar to those on the MNIST (IID) dataset in Fig. 9, showing that adapting the maximum number of LDPC decoding iterations can achieve similar or even slightly better accuracy compared to the baseline scheme. This observation was also reported in [49, 50, 51], which suggests that adding annealed noise to the DNN model could enhance its generalization ability. Moreover, the proposed scheme reduces the LDPC decoding energy consumption on the Fashion-MNIST (IID), CIFAR-10 (IID), and Fashion-MNIST (non-IID) datasets by 18.3%, 20.1%, and 18.0%, respectively. Table III provides a summary of the experimental data.

### v3

@@ III-C LDPC Decoding Iteration Control
$Q_{r}=\frac{1}{\chi_{0}}\mathcal{O}\left(2\log\ (r+1)\right)+\frac{1}{\chi_{0}}\mathcal{O}\left(\frac{1}{\epsilon}\log\frac{1}{\epsilon}\right),$ (17)
which provides a theoretical guideline to determine the maximum number of LDPC decoding iterations in the $r$-th communication round. Specifically, $Q_{r}$ can be set as $\alpha\log(r+1)+\beta$, where $\alpha$ and $\beta$ are constants. However, (17) cannot be directly utilized because $\alpha$ and $\beta$ are code-specific and depend on the receive SNR as well as channel decoding algorithms. Thus, we propose a practical approach via the mapping between BER and the maximum number of LDPC decoding iterations at different levels of receive SNR. Entries in the mapping can be obtained through empirical measurements specific to the LDPC code used in the wireless FL system. Fig. 5 shows the sample simulation results for the rate 1/2 LDPC code with a block length of 1008 bits and binary phase shift keying (BPSK) modulation over an additive white Gaussian noise (AWGN) channel with SNR = 1.5 and 2.5 dB, which can be stored in memory of clients [37]. From the mapping, we are able to obtain the corresponding maximum number of decoding iteration for a target BER for clients with different receive SNR in each communication round.
Figure 5: Mappings between BER and the maximum number of LDPC decoding iterations at different receive SNRs.
The wireless FL procedures with adaptive LDPC decoding are summarized in Algorithm 1. The algorithm starts with an initial BER, denoted as $b_{0}$, and gradually decreases the BER with communication rounds at the rate of $\frac{1}{(r+1)^{2}}$. In practice, since the BER of LDPC decoding cannot be arbitrarily small (see Fig. 5), we set a target BER at the last communication round, denoted as $b_{R-1}$. With $b_{0}$, $b_{R-1}$ ($b_{0}>b_{R-1}$), and the desired BER should scale at the rate of $\frac{1}{(r+1)^{2}}$, $b_{r}$ can be designed as
$b_{r}=\frac{(b_{0}-b_{R-1})R^{2}}{(R^{2}-1)(r+1)^{2}}+\frac{b_{R-1}R^{2}-b_{0}}{R^{2}-1}$ (18)
for $R>1$. The corresponding maximum number of LDPC decoding iterations in that round is obtained from the measured mapping relationship.
@@
Note that the number of entries in the mappings depend on the granularity of discrete BER and SNR values. Since this number is not large, the overhead for storing them is expected to be small. Additionally, reading these entries from memory takes constant time and does not introduce additional LDPC decoding latency.
@@ IV Experimental Results
In this section, we validate the effectiveness of the proposed adaptive LDPC decoding algorithm via numerical experiments. First, we introduce the experimental settings. Then, performance of the proposed algorithm, including test accuracy and decoding energy consumption, is compared with a baseline by fixing the maximum number of LDPC decoding iterations in each communication round (hereinafter referred to as the “Fixed-$Q$” scheme).
Figure 6: BER of a sample client achieved by the baseline scheme (“Fixed-$Q$”) with a fixed maximum number of LDPC decoding iterations: By setting the maximum number of LDPC decoding iterations $Q$ to 6, 12, 18, 24, and 52, clients achieve the BER at around $10^{-1}$, $10^{-2}$, $10^{-3}$, $10^{-4}$, and $10^{-5}$, respectively.
@@ IV-A Experimental Settings
(a) LeNet-300-100 on MNIST (IID).
(b) Vanilla-CNN on Fashion-MNIST (IID).
(c) 7-layers CNN on CIFAR-10 (IID).
(d) Vanilla-CNN on Fashion-MNIST (non-IID).
Figure 7: Test accuracy on different datasets under the baseline scheme “Fixed-$Q$” with a fixed maximum number of LDPC decoding iterations.
(a) Actual number of LDPC decoding iterations.
(b) Distribution of the actual number of LDPC decoding iterations.
Figure 8: Actual numbers of LDPC decoding iterations and its distribution at a sample client in a selected communication round on Fashion-MNIST (IID). A 8-bit vanilla CNN model is encoded by the rate 1/2 min-sum LDPC code of block-length of 1008 bit, resulting in 4620 frames in a communication round.
@@
System Setup: We consider a wireless FL system with 10 clients. The clients communicate with the server via the rate 1/2 min-sum LDPC code using a code length of 1008 bits[38]. BPSK is adopted as the modulation scheme. In each communication round, $N=8$ bits are used to digitalize each model parameter. The number of local training steps in each communication round is set to $E=5$ with the learning rate and batch size as 0.01 and 64, respectively. The receive SNR of all the downlink wireless channels is assumed to be 2.5 dB.
@@
Datasets: We conduct experiments on three classic datasets, including MNIST[39], Fashion-MNIST[40], and CIFAR-10[41]. The training datasets are distributed among clients in an IID manner. We further consider one experiment where the Fashion-MNIST data are non-IID among clients. In particular, each client is assigned with two categories of training data similar to the setup in [42]. The test datasets are used for model accuracy evaluation at the server.
@@
DNN Models: For the MNIST dataset, the LeNet-300-100 model[43] is employed, which is a fully connected neural network consisting of two hidden layers and one softmax layer, with approximately 0.27 million parameters. For the Fashion-MNIST dataset, a vanilla convolutional neural network (CNN) [4] with two 5×5 convolution layers, one fully connected layer, and one softmax layer is employed, comprising around 0.58 million parameters. For the CIFAR-10 dataset, a 7-layer CNN model with four 3×3 convolution layers, two fully connected layers, and one softmax layer is employed, which contains approximately 9.07 million parameters.
IV-B Learning Performance of the Baseline Scheme
(a) $Q_{r}$.
(b) Test accuracy.
(c) Avg. actual number of decoding iterations among clients.
(d) BER at every communication round,
i.e., ($b_{r}$).
(e) Test accuracy at the last communication round.
(f) Total decoding energy consumption.
@@ IV-B Learning Performance of the Baseline Scheme
Figure 9: Experimental results on MNIST (IID). Compared with the baseline scheme, our proposed algorithm that adaptively increases the maximum number of LDPC decoding iterations to 24 after 50 communication rounds can achieve the same final test accuracy of 98% and reduce the decoding energy by 6.7 mJ.
(a) $Q_{r}$.
(b) Test accuracy.
(c) Avg. actual number of decoding iterations among clients.
(d) BER at every communication round,
i.e., ($b_{r}$).
(e) Test accuracy at the last communication round.
(f) Total decoding energy consumption.
@@
Figure 10: Experimental results on Fashion-MNIST (IID). Compared with the baseline scheme, our proposed algorithm that adaptively increases the maximum number of LDPC decoding iterations to 24 after 50 communication rounds can achieve a similar final test accuracy of 91% and reduce the decoding energy by 14 mJ.
@@
In this subsection, experimental results for the baseline scheme “Fixed-$Q$” are presented. First, as shown in Fig. 6, a fixed maximum number of LDPC decoding iterations leads to approximately constant BERs over the training process.
Next, we investigate the impact of a roughly constant BER on the learning performance of wireless FL. Fig. 7 summarizes the test accuracy with different maximum numbers of LDPC decoding iterations. It can be observed that for all the four FL tasks, setting the maximum number of decoding iterations to 6 (which corresponds to a BER of $\sim 10^{-1}$) severely affects the learning accuracy. Similarly, when the maximum number of decoding iterations is set to 12 (which corresponds to a BER of $\sim 10^{-2}$), the training process fails to converge to satisfactory accuracy. By setting the maximum number of decoding iterations as 18 (which corresponds to a BER of $\sim 10^{-3}$), the training process converges, but with a noticeable loss in test accuracy compared to the case when the maximum number of decoding iterations is 24 (which corresponds to a BER of $\sim 10^{-4}$). However, further increasing the maximum number of decoding iterations to 52 does not lead to much accuracy improvement. Thus, we next focus on the most efficient configuration of the baseline scheme with $Q=24$, i.e., “Fixed-24”.
@@ IV-C Learning Performance of the Proposed Scheme
(a) $Q_{r}$.
(b) Test accuracy.
(c) Avg. actual number of decoding iterations among clients.
(d) BER at every communication round,
i.e., ($b_{r}$).
(e) Test accuracy at the last communication round.
(f) Total decoding energy consumption.
@@
Figure 11: Experimental results on CIFAR-10 (IID). Compared with the baseline scheme, our proposed algorithm that adaptively increases the maximum number of LDPC decoding iterations to 24 after 40 communication rounds can achieve a similar final test accuracy of 77.6% and reduce the decoding energy by 191 mJ.
(a) $Q_{r}$.
(b) Test accuracy.
(c) Avg. actual number of decoding iterations among clients.
(d) BER at every communication round,
i.e., ($b_{r}$).
(e) Test accuracy at the last communication round.
(f) Total decoding energy consumption.
@@
Figure 12: Experimental results on Fashion-MNIST (non-IID). Compared with the baseline scheme, our proposed algorithm that adaptively increases the maximum number of LDPC decoding iterations to 24 after 50 communication rounds can achieve a similar final test accuracy of 54.1% and reduce the decoding energy by 13.7 mJ.
@@
In this subsection, the proposed adaptive LDPC decoding scheme is compared to the baseline scheme. Specifically, the learning process of the proposed scheme is assigned with an initial BER of $b_{0}=10^{-1}$ and a final BER of $b_{R-1}=10^{-4}$. The maximum number of LDPC decoding iterations is determined according to its mapping with BER as stated in Section III. In the experiments, the initial maximum numbers of LDPC decoding iterations are thus set to 6 and gradually increased to 24 at the last communication round. For the baseline “Fixed-24” scheme, a maximum of 24 LDPC decoding iterations can be used in each communication round, which corresponds to a BER of $10^{-4}$.
@@
Fig. 8(a) shows an example on the Fashion-MNIST (IID) dataset to illustrate the actual number of decoding iterations of digitalized model parameters at a sample client in a selected communication round. It can observed that only a few frames exhaust all the admissible LDPC decoding iterations, which is set to 52. Note that the decoding energy consumption is proportional to the actual number of decoding iterations. From the distribution of the actual number of LDPC decoding iterations shown in Fig. 8(b), it is observed that 18 decoding iterations are sufficient for 80% of the frames, implying the significant reduction of decoding energy consumption.
@@
Fig. 9 summarizes the experimental results on the MNIST (IID) dataset. Specifically, Fig. 9(a) shows how the maximum number of LDPC decoding iterations ($Q_{r}$) varies with communication rounds in different schemes. While the baseline scheme “Fixed-24” results in a constant BER, the proposed scheme achieves the desired BER by linearly increasing the maximum number of LDPC decoding iterations. We note that this behavior results from the empirical mapping between $Q_{r}$ and $b_{r}$ for the used LDPC decoder, which is different from the general theoretical guideline in (17). Fig. 9(b) shows the test accuracy achieved by the proposed scheme and the baseline “Fixed-24” scheme. Compared with the baseline scheme with a fixed number of 24 decoding iterations, the proposed scheme achieves the same test accuracy with 50 communication rounds despite with slower convergence. Fig. 9(c) illustrates how the average actual number of decoding iterations at clients over all DNN parameters varies during the training process. The proposed scheme can reduce the average number of decoding iterations and thus the overall decoding energy consumption. Fig. 9(d) depicts the BER (i.e., $b_{r}$) of the proposed scheme over communication rounds, which scales as $\mathcal{O}(1/(r+1)^{2})$ from $10^{-1}$ to $10^{-4}$. Fig. 9(e) shows the final test accuracy of different schemes at the 49-th communication round. It can be observed that the proposed scheme achieves the same accuracy as the baseline schemes (“Fixed-24”) that always maintains a very low BER of $10^{-4}$ throughout the training process. Fig. 9(f) illustrates the overall decoding energy consumption at clients under different schemes, which is obtained by multiplying the total number of decoding iterations with the LDPC decoding energy efficiency reported in [23]. Compared to the baseline scheme, the proposed scheme significantly reduces the overall decoding energy consumption by 19.4%.
@@
TABLE III: Summary of Experimental Results. Our scheme can achieve similar accuracy while saving around 20% of the overall decoding energy consumption compared to the best baseline (“Fixed-$Q$”) scheme.
Final Test Accuracy
@@
MNIST (IID) 96.74% 97.96% 98.06% 98.02%
Fashion-MNIST (IID) 86.18% 90.63% 91.24% 91.16%
CIFAR-10 (IID) 76.72% 77.38% 77.57% 77.68%
Fashion-MNIST (non-IID) 51.3% 54.1% 54.3% 54.2%
Total Decoding Energy Consumption (mJ)
@@
MNIST (IID) 33.87 34.51 34.68 27.81
Fashion-MNIST (IID) 74.23 76.10 76.29 62.17
CIFAR-10 (IID) 925.10 948.44 951.36 757.80
Fashion-MNIST (non-IID) 74.22 76.11 76.27 62.38
Figs. 10, 11, and 12 present the experimental results on Fashion-MNIST (IID), CIFAR-10 (IID), and Fashion-MNIST (non-IID) datasets, respectively. The results are similar to those on the MNIST (IID) dataset in Fig. 9, showing that adapting the maximum number of LDPC decoding iterations can achieve comparable or even slightly better accuracy compared to the baseline scheme. This observation was also reported in [44, 45, 46], which suggests that adding annealed noise to the DNN model could enhance its generalization ability. Moreover, the proposed scheme reduces the LDPC decoding energy consumption on the Fashion-MNIST (IID), CIFAR-10 (IID), and Fashion-MNIST (non-IID) datasets by 18.3%, 20.1%, and 18.0%, respectively. Table III provides a summary of the experimental results.